the product of the digits of two digit number is 18. The number formed by interchanging the digits of that number will be 27 more than the original number. Find the original number.
Answer by stanbon(75887)

the product of the digits of two digit number is 18. The number formed by interchanging the digits of that number will be 27 more than the original number. Find the original number.
----
Let the original number be 10t+u
----
Equation:
t*u = 18
10u+t - (10t+u) = 27
----
Simplify:
t*u = 18
9u -9t = 27
----
Modify:
t*u = 18
u - t = 3
-----
Substitute for u and solve for "t":
t(t+3) = 18
t^2 + 3t-18 = 0
(t+6)(t-3) = 0
-----
Positive solution:
t = 3
---
Solve for "u":
tu = 18
u = 6
----
Original Number: 36
========================
Cheers,
Stan H.